There is only one supplier in the market for widgets, which acts as a monopolist. Suppose the monopolist has marginal costs given by MC = 30+QMC = 30+Q. Demand for widgets is given by P = 210−QP = 210−Q.

If the monopolist is maximizing their profits, they will choose to produce a quantity of ____ and charge a price of____.

Points possible: 1

Question 2

What is the deadweight loss caused by the firm acting as a monopolist instead of setting price equal to marginal cost?
Answer #1

Q1

Answer

It produces at MR=MC

MR=210-2Q ....... An MR curve is double sloped than an inverse linear demand curve

MC=30+Q

equating both

210-2Q=30+Q

3Q=180

Q=60

P=210-60=$150

the quantity is 60 units and the price is $150

MC=30+60=90

========

Q2

Answer

DWL =the decrease in the total surplus because of less output and higher price than socially optimum

socially optimum is at MC=P

30+Q=210-Q

2Q=180

Q=90

DWL=0.5*(the price of monopoly -MC of monopoly)*(socially efficient quantity -monopoly quantity)

=0.5*(150-90)*(90-60)

=900

the DWL is $900